OpenStudy (ledah):

Multiplying Radicals: sqrt12 * sqrt3xy^2

OpenStudy (ash2326):

Is this your question? \[\sqrt {12} \times \sqrt{3xy^2}\]

OpenStudy (ledah):

yeah I just don't know how to do the radical thing

OpenStudy (ash2326):

You can use equation editor for this

OpenStudy (ash2326):

Ok let's work on the problem! if we have \[\sqrt{A} \times \sqrt{B}\] if A and B are positive then it's equivalent to \[\sqrt{AB}\] Could you tell what we'll get after multiplying the two radicals?

OpenStudy (ledah):

are we multiplying 12 and 3?

OpenStudy (ash2326):

We are multiplying \[12\times 3 xy^2\]

OpenStudy (ledah):

that leaves 36xy^2

OpenStudy (ash2326):

good, so we have now \[\sqrt{36xy^2}\] Now can you simplify it further?

OpenStudy (ledah):

Not sure. I never learned this in school and my teacher for virtual school is off for the next few days.

OpenStudy (ash2326):

Okay do you know the square root of 36 ?

OpenStudy (ledah):

6

OpenStudy (ash2326):

It's because we \[\sqrt{36}=\sqrt {6\times 6}=6\] here we have \[\sqrt{36xy^2}=\sqrt{6\times 6\times x\times y\times y}\] Could you simplify it now? only the terms appearing twice come out, as 1 term

OpenStudy (ledah):

it would be 6x?

OpenStudy (ash2326):

only y and 6 appear twice, they will come out. x will remain inside \[\sqrt{6\times 6 \times x\times y\times y}=6y\sqrt{x}\] Do you get this?

OpenStudy (ledah):

Oh I didn't quite understand what you meant by "come out" but yes now I understand it.

OpenStudy (ash2326):

Did you understand? What's the square root of 64x^2? \[\sqrt{64x^2}\]

OpenStudy (ledah):

8x?

OpenStudy (ash2326):

good work:D